please help me solve this problem.
Solve the system:
X+y=-5
2x-y=-7

Answer by stanbon(75887)

please help me solve this problem.
Solve the system:
X+y=-5
2x-y=-7
-------------
Add the two equations to get:
3x = -12
x = -4
-----------
Substitute into x+y = -5 to solve for "y":
-4 + y = -5
y = -1
--------------
Cheers,
Stan H.